OpenStudy (anonymous):

Angie and Kenny play online video games. Andie buys 1 software package and 3 months of game play. Kenny buys 1 software package and 2 months of game play. Each software package costs $20. if their total coast is $115, what is the cost of one month of game play? What equation can I use to solve this problem?

OpenStudy (anonymous):

they spent 40 dollars on the softwares

OpenStudy (anonymous):

so 115-40=75

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

75 is how much they have left

OpenStudy (anonymous):

so 75 divided by 5 months of game play

OpenStudy (anonymous):

equals...

OpenStudy (anonymous):

15

OpenStudy (anonymous):

So each month is worth 15 dollars

OpenStudy (anonymous):

now make the equation

OpenStudy (anonymous):

So the equation would be: 115 - 40 = 75 ÷ 5 = x

OpenStudy (anonymous):

you got most of it right

OpenStudy (anonymous):

What part is wrong?

OpenStudy (anonymous):

You need to add the parts where you added the softwares together and the part when you added the months of gameplay together

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

by any chance are you in 6th grade?

OpenStudy (anonymous):

no. lol why?

OpenStudy (anonymous):

because I'm in 6th grade and i did alot of these problems at the beginning of the school year

OpenStudy (anonymous):

no... 10th grade. i feel stupid lol

OpenStudy (anonymous):

oh LOL

OpenStudy (anonymous):

I take advanced classes so....